OpenStudy (anonymous):

Using substitution, determine if x = 120 is a possible solution to the inequality 23 – x ≤ 2x + 8

OpenStudy (anonymous):

yes no

OpenStudy (texaschic101):

try subbing in 120 for x 23 - 120 <= 2(120) + 8 -97 <= 240 + 8 -97 <= 248 yes it is a solution